OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (6, 2) and (4, 8)? y=___x + ___

OpenStudy (susanna):

first, find the slope

OpenStudy (anonymous):

....sorry brain fart. lol how do i do that again?

OpenStudy (er.mohd.amir):

slope=(y2-y1)/(x2-x1)

OpenStudy (astrophysics):

Well first find the slope, using the following formula: \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] where your coordinates represent \[(x_1,y_1)~~\text{and}~~(x_2,y_2)\] Once you manage to get that, plug in m and your first coordinates in the point - slope form: \[y-y _{1}=m(x-x _{1})\] do some little algebra and eventually you will have slope - intercept form \[y=mx+b\]

OpenStudy (susanna):

(8-2)/(4-6)

OpenStudy (susanna):

slope is -3

OpenStudy (anonymous):

i got that too

OpenStudy (susanna):

now use one of the points and the slope. y-y=m(x-x)

OpenStudy (susanna):

(6, 2) and slope =-3 y-2=-3(x-6)

OpenStudy (susanna):

solve for y

OpenStudy (susanna):

y-2=-3x+18

OpenStudy (anonymous):

-3x+ 20

OpenStudy (susanna):

yes y=-3x+20

OpenStudy (anonymous):

yay